Q.5

Becky Shelton, a teacher at Kemp Middle School, is in charge of ordering the T-shirts to be sold for the school’s annual fund-raising project. The T-shirts are printed with a special Kemp School logo. In some years, the supply of T-shirts has been insufficient to satisfy the number of sales orders. In other years, T-shirts have been left over. Excess T-shirts are normally donated to some charitable organization. T-shirts cost the school $7 each and are normally sold for $14 each. Ms. Shelton has decided to order 830 shirts.

Required

  1. If the school receives actual sales orders for 775 shirts, what amount of profit will the school earn? What is the cost of waste due to excess inventory?

  2. If the school receives actual sales orders for 865 shirts, what amount of profit will the school earn? What amount of opportunity cost will the school incur?

$ 385 Profit Waste due to excess inventory Profit Opportunity cost
Answer #1

ANSWER:

Solution (a) :

Cost of shirt purchased = 830 * $7 = $5,810

Actual sales of shirts = 775 * $14 = $10,850

Amount of profit that school earn = sales value - Cost = $10,850- $5,810= $5,040

Cost of waste due to excess inventory = (830 -775 ) * $7 = $385

Solution (b) :

Actual sales order = 865

Actual sales units = Nos of units available i.e. 830 units

Sales value of shirts = 830 *$14 = $11,620

Amount of profit that school will earn = $11,620 - $5,810= $5,810

Amount of opportunity cost that school incur = loss of profit on 100 units

= ($14 - $7) * 100 = $700

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************
Answer #2

Hi. How did you decide to calculate the loss on 100 units? Wouldn't it be the loss of 35 shirts?

answered by: melissa